Project

General

Profile

11222013 Minutes

Attending: Micheal, Gavin, Kanika, Evan, Siva, Mat Jon

- Neutrino scanning work continues.

Operation:
- Cooling work yesterday (6 APD are cold that didn’t before)
- SS HS expected the week after TG
- Trigger: Evan- Broadcast to 5 partitions. This is turned down now to 2 partitions and 3 retries. The retry band is thinner now and evan reports 100% trigger success.
Minos atomic clock questions are out (PA).

Jon:
- Run history services could take minutes to pull thresholds
- Gavin asks about DB tag to match release. Jon - it needs to be implemented
- EVD sometime it’s showing masks that don’t make sense...